OpenStudy (anonymous):

How do you solve the value of m in m^-5 = 1/32

OpenStudy (anonymous):

\[m ^{-5}=\frac{ 1 }{ 32}\]

OpenStudy (dtan5457):

Couple things to remember here \[m^{-5}=\frac{ 1 }{ m^5 }\] \[\frac{ 1 }{ m^5 }=\frac{ 1 }{ 32 }=m^5=32\] \[m=\sqrt[5]{32}\]

OpenStudy (dtan5457):

Now just find the 5th root of 32

OpenStudy (anonymous):

Is 6.4 the right answer?

OpenStudy (dtan5457):

Hmm (6.4)^5=32? I don't think so

OpenStudy (anonymous):

Im not getting it XD is there an other way to explain it??

OpenStudy (jdoe0001):

got calculator? to get the 5th root, no?

OpenStudy (dtan5457):

Think about it the fifth root of 32 implies (x)(x)(x)(x)(x) one number multiplied by itself 5 times will be 32

OpenStudy (dtan5457):

the number in this case is rather small...try 1,2,3..

OpenStudy (anonymous):

wouldnt that be 6.4?? 6.4 times 5 is 32

OpenStudy (dtan5457):

(6.4)x(6.4)x(6.4)x(6.4)x(6.4) does it really?

OpenStudy (anonymous):

OOOOOOH

OpenStudy (anonymous):

I forgot how to get the answer XD How get the root again?? Im sorry

OpenStudy (dtan5457):

What's \[\sqrt{9}\]

OpenStudy (anonymous):

3?

OpenStudy (dtan5457):

But why? Square root of 9 implies What multiplies itself twice is 9, which is three.

OpenStudy (dtan5457):

\[\sqrt[5]{32}\] Similar concept here....what multiplied by itself 5 times equals 32?

OpenStudy (anonymous):

5.656 and so on..?

OpenStudy (anonymous):

Oh wait nevermind

OpenStudy (dtan5457):

You got..?

OpenStudy (anonymous):

Ok i think the problem is that i completely forgot how to get the root. I think thats an issue

OpenStudy (dtan5457):

Calculator..guess and check..

OpenStudy (anonymous):

ok this might take a second XD thank you for trying to help though XD

OpenStudy (anonymous):

OH ITS 2

OpenStudy (dtan5457):

yep. good job

OpenStudy (anonymous):

XD thanks!

OpenStudy (dtan5457):

feel free to fan and medal :) good effort here :)